OpenStudy (unklerhaukus):

\[J = \int\limits_0^\infty\frac{x^3}{e^x-1}\mathrm dx\] Show that \(J = \pi^4/15\).

OpenStudy (zarkon):

you can use \[\frac{1}{e^{x}-1}=\frac{e^{-x}}{1-e^{-x}}=\sum_{k=1}^{\infty}e^{-kx}\] ... convert the problem into finding \(\zeta(4)\)

OpenStudy (unklerhaukus):

... where \(\zeta\) is the Riemann zeta function

OpenStudy (kainui):

I was thinking I could find the poles and turn this into a contour integral, but I haven't had enough practice with them to know how to do this. There's the clear way in which \(e^{ix}-1=0\) has solutions for when \(x=2 \pi n\) except at the origin it looks like it isn't a pole since by L'Hopital's rule it's 0 there. I then tried to look at when \( \frac{e^z-1}{z^3} = 0\) to find the poles to see if there was some kinda simple way to get the residue(s) by turning it into a Laurent series cause I really don't know what I'm doing lol. \[\frac{e^z-1}{z^3} = \sum_{n=1}^\infty \frac{z^{n-3}}{n!} = \sum_{k=-2}^\infty \frac{z^k}{(k+3)!}\] Unfortunately this is not the Laurent series of our function so the k=-1 term is not the residue and I'm basically just sorta fooling around. I could go back to my original thing and do a kind of contour integral maybe like this?  Those are the \(2 \pi i n\) poles and maybe let this thing go out to infinity and hope the arc part goes to 0 as it goes to infinity since there's an exponential in the denominator I believe this much is true. I just don't know about this return path and the fact that I'm including infinitely many poles. Oh well I'll keep thinking about this. Maybe I can just do a single contour containing one pole or alternate between poles?

OpenStudy (owen3):

I just wanted to mention that $$\zeta (4) = {\frac {{\pi }^{4}}{90}}$$

OpenStudy (amilapsn):

Can we do this with double or some higher dimensional integral? Like finding \(\int_0^\infty \frac{\sin x}{x}dx\)? If we can, is there a method to think it?